OpenStudy (imtant):

Hello,everybody.anyone know how to solve this question? Given function of f(x)=I 1-x^2 I for sketch the function for x>-4 & from the graph, find Domain and Range..Anybody?

OpenStudy (campbell_st):

is the question \[f(x) = 1 - x^2\] or \[f(x) = 11 - x^2\]

OpenStudy (imtant):

it is modulus..

OpenStudy (imtant):

yes, f(x)=in modulus 1-x^2..

OpenStudy (anonymous):

\[f(x)=| 1-x^2 | \]

OpenStudy (campbell_st):

ok so its absolute value use a table of values is an easy way  then graph it